# Operator Deep Smoothing for Implied Volatility

### v2

@@ D.3 Example Plots: Option Metrics End-Of-Day US Index Options Data
Figure 16: Smoothing of RUT end-of-day data from 07.01.2021. Every datapoint displayed.
@@ Appendix E SVI
SVI, originally devised by Gatheral (Gatheral,, 2004), stands for Stochastic Volatility Inspired and is a low-dimensional parametrization for implied volatility slices (namely for each maturity). It captures the key features of implied volatility of Equity indices, and has become the industry-wide benchmark for implied volatility smoothing on such markets. Its ”raw” variant parametrizes the ”slice” of implied volatility at a given time-to-expiry $\tau$ as follows:
$\hat{v}_{\tau}(k)=\sqrt{\frac{a+b\left(\rho(k-m)+\sqrt{(k-m)^{2}+\sigma^{2}}\right)}{\tau}},\quad\text{for all }k\in\mathbb{R},$
where $a,b,k,m,\sigma$ are real parameter values.
@@
While stylistically accurate, SVI does not easily guarantee absence of static arbitrage opportunities, and several authors have investigated this issue (Gatheral and Jacquier,, 2014; Martini and Mingone,, 2022; Mingone,, 2022; Martini and Mingone,, 2023). To produce our SVI benchmark we therefore rely on the constrained SLSQP optimizer provided by the SciPy scientific computing package for Python, with the mean square error objective, a positivity constraint and the constraint (2) (computed in closed form), and the following parameter bounds:

### v3

@@ D.3 Example Plots: Option Metrics End-Of-Day US Index Options Data
D.4 Example plots: synthetic SSVI data
SSVI is a continuous interpolation scheme for SVI (see Section E) introduced in Gatheral & Jacquier (2014). We use to generate synthetic volatility surface data, to which we then apply our smoothing method. We consider the following formulation of SSVI (taken from Lind & Gatheral (2023)):
$\hat{v}^{\phi}_{\text{SSVI}}(\tau,k)=\sqrt{\frac{\theta_{\tau}}{2\tau}\left(1+\rho\varphi(\theta_{\tau})k+\sqrt{\left(\varphi(\theta_{\tau})k+\rho\right)^{2}+(1-\rho^{2})}\right)},$
where $\varphi(\theta_{\tau})=\eta\theta_{\tau}^{\gamma}$ and
$\theta_{\tau}=\theta\,\tau+(V-\theta)\frac{1-e^{-\kappa_{1}T}}{\kappa_{1}}+(V^{\prime}-\theta)\frac{\kappa_{1}}{\kappa_{1}-\kappa_{2}}\left(\frac{1-e^{-\kappa_{2}T}}{\kappa_{2}}-\frac{1-e^{-\kappa_{1}T}}{\kappa_{1}}\right),$
making $\phi=(V,V^{\prime},\theta,\rho,p,\eta,\gamma,\kappa_{1},\kappa_{2})$ the parameter vector. We set the parameters to the following values, reported as historical averages for the S&P 500 over the period 1996-2021 in Lind & Gatheral (2023):
$V$ $V^{\prime}$ $\theta$ $\rho$ $p$ $\eta$ $\gamma$ $\kappa_{1}$ $\kappa_{2}$
$0.04$ $0.04$ $0.11$ $-0.5$ $0.01$ $1.19$ $0.49$ $5.5$ $0.1$
We produce synthetic market data $\mathbf{v}_{\text{SSVI}}$ by evaluating the SSVI instance $\hat{v}^{\phi}_{\text{SSVI}}$ on the synthetic $(8\times 51)$-grid
$\pi_{\text{in}}=\pi_{\rho}\times\pi_{z}$
with $\pi_{\rho}=\{0.16,0.28,0.4,0.52,0.64,0.76,0.88,1\}$ and $\pi_{z}=\{-1.5,-1.46,\dots,0.46,0.5\}$. We then proceed to smooth $\mathbf{v}_{\text{SSVI}}$ with our method, producing $\hat{v}_{\text{OpDS}}=\tilde{F}^{\theta}(\mathbf{v}_{\text{SSVI}})$, and plot the results in Figure 19. Moreover, we compute the MAPE smoothing error, both over the grid $\pi_{\text{in}}$, as well as over a high-resolution $(100\times 100)$-grid
$\pi_{\text{out}}=\pi_{\rho}\times\pi_{z},$
where $\pi_{\rho}$ and $\pi_{z}$ discretize $[0.01,1]$ and $[-1.5,0.5]$, respectively, uniformly with one hundred nodes each. We achieve the following values:
$\pi_{\text{in}}$ $\pi_{\text{out}}$
2.01% 2.42%
Figure 19: Smoothing of synthetic SSVI data. Every second datapoint displayed.
@@ Appendix E SVI
SVI, originally devised by Gatheral (Gatheral, 2004), stands for Stochastic Volatility Inspired and is a low-dimensional parametrization for implied volatility slices (namely for each maturity). It captures the key features of implied volatility of Equity indices, and has become the industry-wide benchmark for implied volatility smoothing on such markets. Its raw variant parametrizes the “slice” of implied volatility at a given time-to-expiry $\tau$ as
$\hat{v}^{\phi}_{\tau}(k)=\sqrt{\frac{a+b\left(\rho(k-m)+\sqrt{(k-m)^{2}+\sigma^{2}}\right)}{\tau}},\quad k\in\mathbb{R},$
where $\phi=(a,b,\rho,m,\sigma)$ is the parameter vector.
@@
While stylistically accurate, SVI does not easily guarantee absence of static arbitrage opportunities, and several authors have investigated this issue (Gatheral & Jacquier, 2014; Martini & Mingone, 2022; Mingone, 2022; Martini & Mingone, 2023). To produce our SVI benchmark we therefore rely on the constrained SLSQP optimizer provided by the SciPy scientific computing package for Python, with the mean square error objective, a positivity constraint and the constraint equation 2 (computed in closed form), and the following parameter bounds: